# BiasBusters: Uncovering and Mitigating Tool Selection Bias in Large Language Models

### v1

@@ Abstract
Agents backed by large language models (LLMs) often rely on external tools drawn from marketplaces where multiple providers offer functionally equivalent options. This raises a critical point concerning fairness: if selection is systematically biased, it can degrade user experience and distort competition by privileging some providers over others. We introduce a benchmark of diverse tool categories, each containing multiple functionally equivalent tools, to evaluate tool-selection bias. Using this benchmark, we test seven models and show that unfairness exists with models either fixating on a single provider or disproportionately preferring earlier-listed tools in context. To investigate the origins of this bias, we conduct controlled experiments examining tool features, metadata (name, description, parameters), and pre-training exposure. We find that: (1) semantic alignment between queries and metadata is the strongest predictor of choice; (2) perturbing descriptions significantly shifts selections; and (3) repeated pre-training exposure to a single endpoint amplifies bias. Finally, we propose a lightweight mitigation that first filters the candidate tools to a relevant subset and then samples uniformly, reducing bias while preserving good task coverage. Our findings highlight tool-selection bias as a key obstacle for the fair deployment of tool-augmented LLMs.
@@ 1 Introduction
Large language models (LLMs) have transformed natural language processing, achieving near-human performance on tasks ranging from code generation to creative writing (Naveed et al., 2024; Luo et al., 2024). Yet LLMs cannot directly act in the world: they cannot query databases, fetch live information, or invoke external services. Additionally, their knowledge remains frozen at training time, leaving them prone to “hallucinations” when asked about events beyond their cutoff (Ji et al., 2023). Augmenting LLMs with external “tools” / APIs addresses these shortcomings by allowing models to delegate specialized functions to dedicated services (Qu et al., 2025). It endows LLMs with the ability to act, a core capability often associated with LLM agents (Chowa et al., 2025). A crucial step within the typical tool-usage pipeline is the multi-stage tool-selection process: given an instruction to the LLM, (i) retrieve a short list of the most relevant candidate tools based on the user query (with, e.g., highest semantic similarity) from a potentially large database of tools, (ii) insert their metadata into the prompt, (iii) have the LLM reason and pick one to solve (one of) the necessary user task(s). However, this process introduces a new challenge: bias (see Figure 1). An LLM may prefer certain tools not for their relevance or accuracy, but because of superficial metadata, i.e., tool names, descriptions, or prompt ordering. Such bias can degrade user experience by repeatedly selecting slow or unreliable services and can, therefore, also inflate operational costs. Additionally, under pay-per-request pricing for tool use, consider the scenario where such biases are systematic across frontier LLMs: if they consistently favor tools from a single provider, this creates major market unfairness, disadvantaging competing providers offering functionally equivalent services (RapidAPI, 2025a; Economize, 2024).
@@
In this paper, we make three key contributions. First, we present a large-scale benchmark for measuring tool selection bias, which we use to conduct the first empirical study of tool‐selection bias in LLMs. We introduce total variation–based metrics to quantify selection imbalances, assemble a novel benchmark of clusters of tools with equivalent functionality, and conduct extensive experiments across multiple models and configurations. These tell us that bias exists to a certain extent for all models tested; models either fixate on a single provider or disproportionately prefer earlier-listed tools in context. Second, after confirming the existence of bias, we investigate the root causes of these biases. We define a rich set of tool-level features and use regression and permutation-importance analyses to identify which factors predict selection. We further perturb metadata fields to test their influence and probe whether continued pre-training on biased data can induce persistent preferences. Our findings show that semantic alignment between query and tool description is the strongest interpretable driver of tool choice, metadata interventions on descriptions reliably steer choices, and biased pre-training amplifies preferences. However, none of these fully explain the behavior. Third, we propose a lightweight mitigation strategy: filtering the supplied tools to a subset of relevant ones and sampling uniformly among them. This approach substantially reduces bias while still ensuring the user task can be solved.
@@
By uncovering, explaining, and addressing tool‐selection bias, our work illuminates a critical blind spot in tool‐augmented LLM research. Our contributions go beyond diagnosis: we provide reproducible resources and a simple mitigation strategy that practitioners can adopt immediately. More broadly, we aim to set a foundation for fairer, more reliable tool-calling systems and a precedent to judge tool-calling LLM applications not only by their accuracy but also by the equity of their interactions with external ecosystems. All resources are available online [^1] All resources are available on the public GitHub repository..
@@ 2 Related Work
Fairness & Bias in LLMs. Bias in large language models emerges from their opaque training on vast web‐scale data, often carrying forward and amplifying human prejudices and reasoning shortcuts (Gallegos et al., 2024). Much of the literature has focused on identifying and mitigating these undesirable behaviors (Schick et al., 2021; Gallegos et al., 2024; Bouchard, 2024). Unwanted behaviors in LLMs manifest as social bias, where stereotypes shape outputs and lead to unequal treatment of protected groups, or as cognitive bias, reflecting human-like shortcuts such as anchoring or confirmation effects (Lou & Sun, 2024; Itzhak et al., 2024; O’Leary, 2025). Measuring these phenomena requires domain‐specific metrics. For example, to measure how likely a model is to associate a neutral attribute with one group over another, one can utilize the normalized log-probability bias score (Kurita et al., 2019). Mitigation strategies are diverse and can be deployed at multiple stages (e.g., at pre-/post-inference via augmentation of training data or rewriting of outputs or during training via fairness-aware objectives) (Gallegos et al., 2024). However, no single approach suffices for every scenario; effective debiasing requires interventions matched to the task’s fairness goals and deployment context. We are not aware of prior work on bias in the tool-use paradigm.
@@
Fairness in Tool Selection. Research on tool-selection bias is limited; however, some works in information retrieval address related issues such as positional or ranking bias (Dai et al., 2024; Ziems et al., 2024). Similarly, in code generation, favoritism for certain providers has been observed  (Zhang et al., 2025). However, these studies do not examine the specific setting of tool selection among providers. Recent studies do show that LLMs can develop systematic preferences in this setting, for example, for selecting tools with manipulated metadata (Mo et al., 2025; Faghih et al., 2025). Analogously, one study shows an advanced attack on tool selection which splits malicious tool entries into fragments optimized for retrieval and selection, achieving high success rates while evading defenses (Shi et al., 2025). These findings reveal that tool-selection mechanisms are highly fragile and susceptible to both naïve and advanced exploitation. In contrast to overt adversarial attacks, our study focuses on subtler sources of bias: small, non-adversarial differences in phrasing or metadata that can still distort fairness in tool selection.
@@ 4.1 Experimental set-up
Models. We evaluate a diverse set of LLMs, including ChatGPT-3.5-turbo and 4.1-mini by OpenAI (2025), Claude 3.5 Sonnet by Anthropic (2025), DeepSeek R1 (DeepSeek-AI et al., 2025), Gemini 2.5 Flash (Google, 2023), ToolLLaMA-2-7B (Qin et al., 2024), and Qwen3 models spanning 1.7B to 235B parameters (Bai et al., 2023).
@@ 4.2 How do LLMs select among functionally equivalent APIs?
To get an idea of how biased each model is in their tool-selection behavior, we compute the positional bias $\delta_{\mathrm{pos}}$, API bias $\delta_{\mathrm{API}}$, and their average $\delta_{\mathrm{model}}$ for seven different models (Section 3.1). Table 1 reports these metrics averaged across all clusters and three runs. All models exhibit substantial bias: $\delta_{\mathrm{model}}$ values are around 0.3–0.4, meaning that roughly 30–40% of the selection probability mass would have to be redistributed to achieve fairness. ChatGPT 4.1-mini is especially biased, with a combined metric of $\sim$0.38. The least biased model in our suite is Qwen 3 (235B), which attains the lowest $\delta_{\mathrm{model}}$.
@@
ChatGPT 3.5-turbo 0.320±.022 0.336±.012 0.328±.005
ChatGPT 4.1-mini 0.331±.008 0.423±.002 0.377±.004
@@
Ablation and Sensitivity Analysis. We also analyze the robustness of tool-selection bias across several factors (full results in Appendix E). Temperature: raising temperature reduces bias modestly by softening extreme preferences. Top-p: increasing top-p has only a negligible effect. Model size: larger models exhibit noticeably less bias than smaller ones. API ordering: cyclic vs. random permutations produce very similar outcomes, indicating intrinsic preferences dominate. System prompts: rewording or restructuring prompts shifts which tools are favored, but does not remove bias.
@@ 4.3 Are tool-selections driven by human-interpretable heuristics?
Figure 5: Mean total-variation (TV) distance from the base selection distribution (no perturbation) to the distribution pertaining to each metadata perturbation (higher = larger shift). Blue bars show results for Gemini and orange bars for ChatGPT. The error bars denote standard deviation across clusters. The run-to-run standard deviation is left out; max run-to-run variability of per-run mean TV was $0.084$.
@@
Figure 5 shows the TV distance from the base selection distribution to the distribution obtained after a certain metadata perturbation, averaged over clusters and experimental runs. It shows a clear hierarchy: logically minimizing semantic signal by scrambling the description, name, and parameters causes the biggest shift. For Gemini, corrupting descriptions+parameters yields the second largest, most reliable shifts in selection (0.450 ± 0.203 TV), followed by scrambling the description of the most popular tool (0.419 ± 0.179), and even description swaps meaningfully steer choices (0.338 ± 0.205). By contrast, name-only perturbations are smaller and have higher variance, and parameters-only are the least impactful. We see similar results for ChatGPT; description perturbation is more impactful than name/parameter perturbations. Overall sensitivity is higher for Gemini than ChatGPT (mean TV 0.310 vs 0.234), indicating greater responsiveness to metadata changes.
@@ 4.4 Can we mitigate the observed bias?
Figure 6 shows the empirical selection distributions of ChatGPT 4.1-mini before and after applying our mitigation method over three clusters of functionally equivalent APIs. Where choices were far from uniform before, after the mitigation method is applied we notice an even spread of selection share. However, in the weather forecasting cluster (right), we see that one API is still being underutilized, perhaps being an indication that the model has a difficult time using this API correctly even when it is the only API available. The effectiveness of the mitigation method is further exemplified in Table 5, where a steep decrease in all our bias metrics after applying the mitigation method can be observed.
@@
Figure 6: Selection distributions for ChatGPT 4.1-mini with and without utilizing our mitigation method across three clusters. Each subplot corresponds to one cluster, with the x-axis indicating the API in the cluster and the y-axis showing the fraction of times that API was chosen using the corresponding setup over 500 inference runs.
@@ 5 Conclusion
Limitations: This study is limited by using 100 synthetically generated user queries per cluster and a narrow seven-feature set for selection-rate predictions, limited model/repeat coverage, and a focus on APIs from RapidAPI and English queries. Although our choices were constrained by compute (our setup already required $\sim$500,000 inference runs), these factors may induce variance and restrict generality. Future Work: Future work should scale queries and clusters and enrich features with deeper semantic and structural signals to raise predictive power beyond the modest observed $R^{2}$. In addition, deploying more expressive models (e.g., boosted trees or deep nets) with cross-validation could capture higher-order interactions between tool features, further increasing their explanatory power. Lastly, broader replication across LLMs and runs would aid in quantifying variability.

### v2

@@ Abstract
Agents backed by large language models (LLMs) increasingly rely on external tools drawn from marketplaces where multiple providers offer functionally equivalent options. This raises a critical fairness concern: systematic bias in tool selection can degrade user experience and distort competition by privileging certain providers over others. We introduce a benchmark of diverse tool categories, each containing multiple functionally equivalent tools, to systematically evaluate tool-selection bias. Using this benchmark, we evaluate seven LLMs and show that substantial bias persists, with models either fixating on a single provider or disproportionately favoring tools that appear earlier in the context. To uncover the sources of this behavior, we conduct controlled experiments that isolate the effects of tool features, exposed metadata (name, description, and parameters), and pre-training exposure. We find that (1) semantic alignment between user queries and tool metadata is the strongest driver of selection; (2) small perturbations to tool descriptions can significantly shift choices; and (3) repeated pre-training exposure to a single endpoint amplifies provider-level bias. Finally, we propose a lightweight mitigation strategy that first filters tools to a relevant subset and then samples uniformly, substantially reducing selection bias while maintaining strong task coverage. Our results highlight tool-selection bias as a key obstacle to the fair deployment of tool-augmented LLM agents. Our code and benchmark are publicly available at https://github.com/thierry123454/tool-selection-bias.
@@ 1 Introduction
Large language models (LLMs) have transformed natural language processing, achieving near-human performance on tasks ranging from code generation to creative writing (Naveed et al., 2024; Luo et al., 2024). Yet LLMs cannot directly act in the world: they cannot query databases, fetch live information, or invoke external services. Additionally, their knowledge remains frozen at training time, leaving them prone to “hallucinations” when asked about events beyond their cutoff (Ji et al., 2023). Augmenting LLMs with external “tools” / APIs addresses these shortcomings by allowing models to delegate specialized functions to dedicated services (Qu et al., 2025). It endows LLMs with the ability to act, a core capability often associated with LLM agents (Chowa et al., 2025). A crucial step within the typical tool-usage pipeline is the multi-stage tool-selection process: given an instruction to the LLM, (i) retrieve a short list of the most relevant candidate tools based on the user query (with, e.g., highest semantic similarity) from a potentially large database of tools, (ii) insert their metadata into the prompt, (iii) have the LLM reason and pick one to solve (one of) the necessary user task(s) (see Appendix K covering the entire tool-usage pipeline). However, this process introduces a new challenge: bias (see Figure 1). An LLM may prefer certain tools not for their relevance or accuracy, but because of superficial metadata, i.e., tool names, descriptions, or prompt ordering. Such bias can degrade user experience by repeatedly selecting slow or unreliable services and can, therefore, also inflate operational costs. Additionally, under pay-per-request pricing for tool use, consider the scenario where such biases are systematic across frontier LLMs: if they consistently favor tools from a single provider, this creates major market unfairness, disadvantaging competing providers offering functionally equivalent services (RapidAPI, 2025a; Economize, 2024). See Appendix L covering the consequences of tool-selection bias in more detail.
In this paper, we make three key contributions.
First, we present a large-scale benchmark for measuring tool selection bias, which we use to conduct the first empirical study of tool‐selection bias in LLMs. We introduce total variation–based metrics to quantify selection imbalances, assemble a novel benchmark of clusters of tools with equivalent functionality, and conduct extensive experiments across multiple models and configurations. These tell us that bias exists to a certain extent for all models tested; models either fixate on a single provider or disproportionately prefer earlier-listed tools in context.
@@
Second, after confirming the existence of bias, we investigate the root causes of these biases. We define a rich set of tool-level features and use regression and permutation-importance analyses to identify which factors predict selection. We further perturb metadata fields to test their influence and probe whether continued pre-training on biased data can induce persistent preferences. Our findings show that semantic alignment between query and tool description is the strongest interpretable driver of tool choice, metadata interventions on descriptions reliably steer choices, and biased pre-training amplifies preferences. However, none of these fully explain the behavior.
Third, we propose a lightweight mitigation strategy: filtering the supplied tools to a subset of relevant ones and sampling uniformly among them. This approach substantially reduces bias while still ensuring the user task can be solved.
@@
By uncovering, explaining, and addressing tool‐selection bias, our work illuminates a critical blind spot in tool‐augmented LLM research. Our contributions go beyond diagnosis: we provide reproducible resources and a simple mitigation strategy that practitioners can adopt immediately. More broadly, we aim to set a foundation for fairer, more reliable tool-calling systems and a precedent to judge tool-calling LLM applications not only by their accuracy but also by the equity of their interactions with external ecosystems.
@@ 2 Related Work
Fairness & Bias in LLMs. Bias in large language models emerges from their opaque training on vast web‐scale data, often carrying forward and amplifying human prejudices and reasoning shortcuts (Gallegos et al., 2024). Much of the literature has focused on identifying and mitigating these undesirable behaviors (Schick et al., 2021; Gallegos et al., 2024; Bouchard, 2024). Unwanted behaviors in LLMs manifest as social bias, where stereotypes shape outputs and lead to unequal treatment of protected groups, or as cognitive bias, reflecting human-like shortcuts such as anchoring or confirmation effects (Lou & Sun, 2024; Itzhak et al., 2024; O’Leary, 2025). Measuring these phenomena requires domain‐specific metrics. For example, to measure how likely a model is to associate a neutral attribute with one group over another, one can utilize the normalized log-probability bias score (Kurita et al., 2019). More recently, formal certification frameworks have been proposed to quantify counterfactual bias over distributions of prompts with statistical guarantees (Chaudhary et al., 2025). Mitigation strategies are diverse and can be deployed at multiple stages (e.g., at pre-/post-inference via augmentation of training data or rewriting of outputs or during training via fairness-aware objectives) (Gallegos et al., 2024). However, no single approach suffices for every scenario; effective debiasing requires interventions matched to the task’s fairness goals and deployment context. We are not aware of prior work on bias in the tool-use paradigm.
Positional Bias in LLMs. Prior research has revealed significant positional selection biases in how LLMs answer multiple-choice questions. Pezeshkpour & Hruschka (2024) and Zheng et al. (2024) demonstrated that LLMs exhibit sensitivity to the ordering of answer options, with performance varying systematically based on option position. This was corroborated in the investigation of Wei et al. (2024), who also found token sensitivity, where specific token characteristics at different positions affect selection. Recently Wang et al. (2025), analyses the underlying model mechanisms that give rise to position bias and proposed a mechanistic approach to eliminating it. In this work, we study and investigate a broad range of potential biases in the critical context of tool selection.
@@
Fairness in Tool Selection. Research on tool-selection bias is limited; however, some works in information retrieval address related issues such as positional or ranking bias (Dai et al., 2024; Ziems et al., 2024). Similarly, in code generation, favoritism for certain providers has been observed  (Zhang et al., 2025b). However, these studies do not examine the specific setting of tool selection among providers. Recent studies do show that LLMs can develop systematic preferences in this setting, for example, for selecting tools with manipulated metadata (Mo et al., 2025; Faghih et al., 2025; Zhang et al., 2025a). Analogously, one study shows an advanced attack on tool selection which splits malicious tool entries into fragments optimized for retrieval and selection, achieving high success rates while evading defenses (Shi et al., 2025). Other recent work has focused on adversarial robustness of tool selection, proposing statistical certification of tool-selection accuracy under adaptive attacks (Yeon et al., 2025). These findings reveal that tool-selection mechanisms are highly fragile and susceptible to both naïve and advanced exploitation. In contrast to overt adversarial attacks, our study focuses on subtler sources of bias: small, non-adversarial differences in phrasing or metadata that can still distort fairness in tool selection.
@@ 4.1 Experimental set-up
Models. We evaluate a diverse set of LLMs, including GPT-3.5-turbo and GPT-4.1-mini by OpenAI (2025), Claude 3.5 Sonnet by Anthropic (2025), DeepSeek-V3.2-Exp (DeepSeek-AI et al., 2025), Gemini 2.5 Flash (Google, 2023), ToolLLaMA-2-7B (Qin et al., 2024), and Qwen3 models spanning 1.7B to 235B parameters (Bai et al., 2023).
@@ 4.2 How do LLMs select among functionally equivalent APIs?
To get an idea of how biased each model is in their tool-selection behavior, we compute the positional bias $\delta_{\mathrm{pos}}$, API bias $\delta_{\mathrm{API}}$, and their average $\delta_{\mathrm{model}}$ for seven different models (Section 3.1). Table 1 reports these metrics averaged across all clusters and three runs. All models exhibit substantial bias: $\delta_{\mathrm{model}}$ values are around 0.3–0.4, meaning that roughly 30–40% of the selection probability mass would have to be redistributed to achieve fairness. GPT-4.1 mini is especially biased, with a combined metric of $\sim$0.38. The least biased model in our suite is Qwen 3 (235B), which attains the lowest $\delta_{\mathrm{model}}$.
@@
GPT-3.5-turbo 0.320±.022 0.336±.012 0.328±.005
GPT-4.1 mini 0.331±.008 0.423±.002 0.377±.004
@@
Ablation and Sensitivity Analysis. We also analyze the robustness of tool-selection bias across several factors (full results in Appendix E). Temperature: raising temperature reduces bias modestly by softening extreme preferences. Top-p: increasing top-p has only a negligible effect. Model size: larger models exhibit noticeably less bias than smaller ones. API ordering: cyclic vs. random permutations produce very similar outcomes, indicating intrinsic preferences dominate. System prompts: rewording or restructuring prompts shifts which tools are favored, but does not remove bias. Toolset size: the composition of the toolset given in context is the primary driver of selection behavior, more than the size of the toolset alone (see Appendix J).
@@ 4.3 Are tool-selections driven by human-interpretable heuristics?
Figure 5: Mean total-variation (TV) distance from the base selection distribution (no perturbation) to the distribution pertaining to each metadata perturbation (higher = larger shift). Blue bars show results for Gemini and orange bars for GPT-3.5. The error bars denote standard deviation across clusters. The run-to-run standard deviation is left out; max run-to-run variability of per-run mean TV was $0.084$.
@@
Figure 5 shows the TV distance from the base selection distribution to the distribution obtained after a certain metadata perturbation, averaged over clusters and experimental runs. It shows a clear hierarchy: logically minimizing semantic signal by scrambling the description, name, and parameters causes the biggest shift. For Gemini, corrupting descriptions+parameters yields the second largest, most reliable shifts in selection (0.450 ± 0.203 TV), followed by scrambling the description of the most popular tool (0.419 ± 0.179), and even description swaps meaningfully steer choices (0.338 ± 0.205). By contrast, name-only perturbations are smaller and have higher variance, and parameters-only are the least impactful. We see similar results for GPT; description perturbation is more impactful than name/parameter perturbations. Overall sensitivity is higher for Gemini than GPT (mean TV 0.310 vs 0.234), indicating greater responsiveness to metadata changes.
@@ 4.4 Can we mitigate the observed bias?
Figure 6 shows the empirical selection distributions of GPT-4.1 mini before and after applying our mitigation method over three clusters of functionally equivalent APIs. Where choices were far from uniform before, after the mitigation method is applied we notice an even spread of selection share. However, in the weather forecasting cluster (right), we see that one API is still being underutilized, perhaps being an indication that the model has a difficult time using this API correctly even when it is the only API available. The effectiveness of the mitigation method is further exemplified in Table 5, where a steep decrease in all our bias metrics after applying the mitigation method can be observed.
@@
Figure 6: Selection distributions for GPT-4.1 mini with and without utilizing our mitigation method across three clusters. Each subplot corresponds to one cluster, with the x-axis indicating the API in the cluster and the y-axis showing the fraction of times that API was chosen using the corresponding setup over 500 inference runs.
@@ 5 Conclusion
Limitations: This study is limited by using 100 synthetically generated user queries per cluster and a narrow seven-feature set for selection-rate predictions, limited model/repeat coverage, and a focus on APIs from RapidAPI and English queries. Although our choices were constrained by compute (our setup already required $\sim$500,000 inference runs), these factors may induce variance and restrict generality.
Future Work: Future work should scale queries and clusters and enrich features with deeper semantic and structural signals to raise predictive power beyond the modest observed $R^{2}$. In addition, deploying more expressive models (e.g., boosted trees or deep nets) with cross-validation could capture higher-order interactions between tool features, further increasing their explanatory power.
We also believe it would be valuable to isolate the contribution of preference tuning (e.g., RLHF) to tool-selection bias, since this might implicitly reward particular writing styles or phrasings in tool names/descriptions and thereby skew selection among otherwise equivalent tools. Finally, richer analysis of models’ internal decision processes could strengthen interpretability: while we now allow a single reasoning step before tool choice, preliminary inspection suggests models often do not explicitly justify why a specific tool is selected. Allowing longer reasoning traces and conducting more systematic chain-of-thought analysis may yield clearer causal explanations of selection behavior, but we leave this for future work due to computational constraints. Lastly, broader replication across LLMs and runs would aid in quantifying variability.
Acknowledgments
JY acknowledges support from Microsoft Ltd. AB and PT acknowledge the 2025 UK AISI Systemic Safety Grant and the UKRI Turing AI Fellowship (EP/W002981/1). AB, JY and PT are also affiliated with the Institute for Decentralized AI, which is supported by an AI Safety Fund grant. This work is supported by the European commissions’ DVPS project. This work is supported by the Schmidt AI 2050.
Ethics Statement
This research did not involve identifiable human data or animals and therefore did not require approval from an institutional ethics committee or review board. All experiments are conducted for scientific purposes only. The work does not involve or target any sensitive attributes such as gender, race, nationality, or skin color. Our study focuses on identifying and minimizing tool-selection bias in LLM agents, with the aim of improving the trustworthiness and safety of the deployment of LLM agents.
Reproducibility Statement
We have made every effort to ensure the reproducibility of our work. We provide detailed descriptions of data and experimental setup in Section 4. Our code is available as a public GitHub repository to facilitate replication.
Additional note. During initial experiments we observed an unusually large standard deviation in DeepSeek’s selection distributions for certain clusters. When we re-ran the experiments under the same configuration, this anomaly disappeared. Since we use the generic deepseek-chat endpoint for our LLM calls, which always serves the latest model version, one of the later runs was executed after a silent model update and thus, ran with a different model. All DeepSeek results reported in the current version of the paper are based on a fresh set of runs collected using the same model, which removes the anomalous error bars.